Blinkeria is considering introducing a new line of hand scanners that can be used to copy material and then download it into a personal computer. These scanners are expected to sell for an average price of $104 each, and the company analysts performing the analysis expect that the firm can sell 108,000 units per year at this price for a period of five​ years, after which time they expect demand for the product to end as a result of new technology. In​ addition, variable costs are expected to be $20 per unit and fixed​ costs, not including​ depreciation, are forecast to be $1,010,000 per year. To manufacture this​ product, Blinkeria will need to buy a computerized production machine for $10.9 million that has no residual or salvage​ value, and will have an expected life of five years. In​ addition, the firm expects it will have to invest an additional $302,000 in working capital to support the new business. Other pertinent information concerning the business venture is provided​ here:

Initial cost of the machine

​$10,900,000

Expected life

5 years

Salvage value of the machine

​$0

Working capital requirement

​$302,000

Depreciation method

straight line

Depreciation expense

​$2,180,000 per year

Cash fixed

costslong dash—excluding

depreciation

​$1,010,000 per year

Variable costs per unit

​$20

Required rate of return or cost of capital

9.2​%

Tax rate

34​%

a. Calculate the​ project's NPV.

b.  Determine the sensitivity of the​ project's NPV to​ a(n) 9 percent decrease in the number of units sold.

c. Determine the sensitivity of the​ project's NPV to​ a(n) 9 percent decrease in the price per unit.

d. Determine the sensitivity of the​ project's NPV to​ a(n) 9 percent increase in the variable cost per unit.

e. Determine the sensitivity of the​ project's NPV to​ a(n) 9 percent increase in the annual fixed operating costs.
Answer #1
a) CALCULATION OF PROJECT NPV
A Unit sales 108000
B Unit Sales Price $104
C=A*B Annual Sales Revenue $11,232,000
D Variable Cost Per unit $20
E=A*D Total Variable Costs $2,160,000
F=C-E Contribution Margin $9,072,000
G Annual Depreciation $2,180,000
H Annual Fixed Cost(excluding Depreciation) $1,010,000
I=F-G-H Income Before Taxes $5,882,000
J Tax Rate 34%
K=I*J Tax Expenses $1,999,880
L=I-K Net Operating Income $3,882,120
M Add: Depreciation (non cash expense) $2,180,000
N=L+M Annual Operating Cash Flow $6,062,120
DISCOUNT FACTOR:
Rate Discount Rate =Required Return 9.2%
Nper Number of Years 5
PV Present Value of annuity of $1                   3.8696 (Using PV function of excel with Rate=9.2%,Nper=5, Pmt=-1)
DF DISCOUNT FACTOR=                   3.8696
P=N*DF Present Value of Annual Cash inflow $23,457,677
PW Present Value of Working Capital Released in year 5 $196,279 (302000/(1.09^5)
PV=P+PW Total Present Value of Positive Cash Flows $23,653,956
IC Initial Cost of machine $10,900,000
IW Initial Working Capital $302,000
I=IC+IW Total Initial Cash Outflow $11,202,000
NPV=PV-I Net Present Value (NPV) $12,451,956
b) NPV with 9% DECREASE in number of units sold
A Unit sales 98280 (108000*(1-0.09)
B Unit Sales Price $104
C=A*B Annual Sales Revenue $10,221,120
D Variable Cost Per unit $20
E=A*D Total Variable Costs $1,965,600
F=C-E Contribution Margin $8,255,520
G Annual Depreciation $2,180,000
H Annual Fixed Cost(excluding Depreciation) $1,010,000
I=F-G-H Income Before Taxes $5,065,520
J Tax Rate 34%
K=I*J Tax Expenses $1,722,277
L=I-K Net Operating Income $3,343,243
M Add: Depreciation (non cash expense) $2,180,000
N=L+M Annual Operating Cash Flow $5,523,243
DISCOUNT FACTOR:
Rate Discount Rate =Required Return 9.2%
Nper Number of Years 5
PV Present Value of annuity of $1                   3.8696
DF DISCOUNT FACTOR=                   3.8696
P=N*DF Present Value of Annual Cash inflow $21,372,466
PW Present Value of Working Capital Released in year 5 $196,279
PV=P+PW Total Present Value of Positive Cash Flows $21,568,745
IC Initial Cost of machine $10,900,000
IW Initial Working Capital $302,000
I=IC+IW Total Initial Cash Outflow $11,202,000
NPV=PV-I Net Present Value (NPV) $10,366,745
H28 X f C D E =PV(H26,H27,-1) F Annual Depreciation Annual Fixed Cost(excluding Depreciation) I=F-G-H Income Before Taxes Tax
C=A*B E=A*D FEC-E 108000 $94.64 (104*(1-0.09) $10,221,120 $20 $2,160,000 $8,061,120 $2,180,000 $1,010,000 $4,871,120 34% $1,6d) C=A*B E=A*D FEC-E G 108000 $104 $11,232,000 $21.80 (20*1.09) $2,354,400 $8,877,600 $2,180,000 $1,010,000 $5,687,600 34% $1C=A*B E=A*D FEC-E I=F-G-H K=1*] L=I-K M NPV with 9% INCREASE in Annual Fixed Operating costs Unit sales 108000 Unit Sales Pri